OpenStudy (anonymous):

7/10n = 1/4 ( what is n) n/2= 4/5 (what is n) n/8= 6/7 (what is n) 1 2/3n = 5/8 (what is n) these are the choices: 1 3/5, 3/8, 5/14, 6 6/7

OpenStudy (anonymous):

Ok hello @SolomonZelman

OpenStudy (solomonzelman):

I am here.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you have the answers for the problems

OpenStudy (anonymous):

you have to find the missing fraction for n

OpenStudy (anonymous):

@SolomonZelman please help fast, I need to submit this

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

hi @angel12310

OpenStudy (solomonzelman):

Is this cool?

OpenStudy (anonymous):

wait LOL

OpenStudy (anonymous):

will this help:

OpenStudy (solomonzelman):

No, ik I am right, just like always, I mean, like you got it?

OpenStudy (anonymous):

i still dont get it, which one is it for 7/10n = 1/4?

OpenStudy (anonymous):

7/10n = 1/4? n=3/8

OpenStudy (solomonzelman):

@angel12310, that's quite cool! But that's not necessary at all.

OpenStudy (anonymous):

yes that is the answer

OpenStudy (anonymous):

OK THANKS @angel12310 AND WHICH ONE IS FOR n/2= 4/5

OpenStudy (anonymous):

n/2 = 4/5 n=6 6/7

OpenStudy (solomonzelman):

5/14 8/5 or 1 3/5 48/7 or 6 6/7 3/8

OpenStudy (anonymous):

n/8= 6/7 n=1 3/5

OpenStudy (anonymous):

n/8= 6/7

OpenStudy (solomonzelman):

I said this 10 times.

OpenStudy (anonymous):

1 2/3n = 5/8 n=5/14

OpenStudy (anonymous):

sorry @SolomonZelman

OpenStudy (anonymous):

thank you @angel12310

OpenStudy (anonymous):

those are all the answers

OpenStudy (anonymous):

@angel12310 i have one more question

OpenStudy (anonymous):

what is it

OpenStudy (solomonzelman):

Are you guys in the same class?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Which of the following shows the right side of the equation after applying a reasonable first step to solve -h/2 + -1/4= 4/9 ? Choices: 4/9 - 1/4, 4/9 + 4/9, 4/9 + 1/4, or 4/9 - 4/9

OpenStudy (anonymous):

don't know this one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know @SolomonZelman

OpenStudy (solomonzelman):